Jackie got a loan at 7% and another at 10% amounting to php 27 000. If the annual interest on both is php 2 250, how much did she borrow at each rate?!
Answer by ikleyn(52781)

.
    x +     y = 27000        (1)    (where x and y are amounts of the loans)
0.07x + 0.10y =  2250        (2)


From eq(1), express x = 27000-y  and substitute it into eq(2), replacing y. You will get


0.07*(27000-y) + 0.10y = 2250.


Solve it for y.  Then evaluate  x = 27000-y.

--------------------

If you need more details,  or if you want to see other similar problems solved by different methods,  look into the lesson
    - Using systems of equations to solve problems on investment
in this site.

You will find there different approaches  (using one equation or a system of two equations in two unknowns),  as well as
different methods of solution to the equations  (Substitution,  Elimination).

Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lesson is the part of this online textbook under the topic  "Systems of two linear equations in two unknowns".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.